OpenStudy (anonymous):

Solve: x2 + 17x = -30

OpenStudy (anonymous):

@Hero and @uri

OpenStudy (anonymous):

factorize it

Directrix (directrix):

You can solve by factoring. x2 + 17x = -30 x^2 + 17x + 30 = 0 To factor, what are two numbers that multiply to 30 AND add to 17?

OpenStudy (anonymous):



OpenStudy (anonymous):

now x^+2x+15x+30

OpenStudy (anonymous):

x(x+2)+15(x+2) (x+2)(x+15) x=-2,-15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

anytime